# Top-of-atmosphere radiation fields over the last millennium reconstructed from proxies

### v1

@@ 5.4 Climate response to volcanic eruptions
Large volcanic eruptions are the dominant pre-industrial climate forcing (Büntgen et al., 2020). Clusters of them are thought to be the main cause of the LIA (Miller et al., 2012; Brönnimann et al., 2019) and previous cold periods (Van Dijk et al., 2024). Sea-surface cooling after eruptions is initially damped by the upper ocean and absorbed into the deeper ocean (Gregory et al., 2016). The cold anomalies then slowly resurface, extending the cooling well beyond the volcanic aerosol residence time of 1–2 years. A decadal pacing of large eruptions would thus allow OHC anomalies to accumulate and persist (Gupta and Marshall, 2018; Zhong et al., 2011). In our volcanic composite, the OHC does not begin to recover until 5–10 years after the eruption (Fig. 10b). If another eruption were to occur during this period, the heat loss would compound. We found that this has happened multiple times over the last millennium, e.g., in the early 1100s, late 1200s, early 1700s, and early 1800s, causing upper-ocean heat loss of 50 ZJ or more over 10 years (Fig. S15). The late 1200s cluster is thought to have initiated the LIA (Miller et al., 2012; Zhong et al., 2011). Indeed, the decade with the lowest EEI over the last millennium is 1283–1292 (0.6 W m-2 below 1850-1900 mean), with a minimum in 1284 (0.8 W m-2 below 1850-1900 mean). However, the start of the cooling trend and a multidecadal cold period align better with the early 1100s cluster, although the energy loss was less sustained. The LIA may thus have been initiated by multiple clusters (both annotated in Fig. 5b).
@@
The upper-ocean heat loss differs qualitatively between our reconstruction and model simulations (Fig. 10 and S14). The simulated ocean cooling peaks 3 years after the eruption and recovers immediately, while the reconstructed cooling peaks at a lower amplitude. In contrast to the simulations, the reconstructed onset of recovery varies by eruption, occurring 3–10 years post-eruption and resulting in the flatter composite (Fig. 10b). This damped and lagged response is potentially a proxy artifact since the biological memory in tree-ring width proxies, which form the bulk of our network, smooths out the volcanic response (Lücke et al., 2019; Zhu et al., 2020). However, models are also known to overestimate the cooling after large-magnitude eruptions when prescribing aerosol optical properties rather than simulating interactive aerosols (Marshall et al., 2025). The discrepancy can be further reduced by accounting for uncertainty in VSSI and timing (Lücke et al., 2023). The timescale of OHC recovery in our reconstruction agrees with instrumental datasets for three large eruptions (Fig. 4).
The loss of OHC is driven by a negative EEI at the TOA, arising from increased RSR due to stratospheric aerosols (Fig. 10c,d). After 5–10 years, the EEI becomes marginally positive and the OHC recovers over the course of a decade. The recovery is associated with an RSR that is slightly less than pre-eruption, rather than reduced OLR, also evident in coupled model simulations (Fig. S14c). The recovery is aided by deep ocean heat release $\gamma T$, which results from the difference between the EEI and dOHC300/dt in Fig. 10d.
There is cooling both after volcanic eruptions and in the last-millennium temperature trend, but the spatial patterns of cooling and radiation differ (Figs. S11 and S13). The millennial trend shows strong polar amplification and a colder equatorial Pacific, which induces a La Niña-like OLR response. In contrast, volcanic eruptions cause more cooling over land and actually warm the equatorial Pacific slightly so that the tropical OLR response is El Niño-like, even though the extratropical OLR is reduced in both cases. The volcanic RSR is positive everywhere due to the global dispersal of stratospheric aerosols. While the post-eruption warming of the equatorial Pacific has been observed before, it is indistinguishable from natural variability. Indeed, our Niño 3.4 index does not change significantly after volcanic eruptions (Fig. S12b), consistent with Zhu et al. (2022) and Dee et al. (2020).
@@
Previous modeling studies have found expanded sea ice persisting for more than a century following large eruptions due to sea ice–ocean feedbacks in the Arctic Ocean and North Atlantic (Zhong et al., 2011; Slawinska and Robock, 2018). In contrast, in our volcanic composite, the sea ice area in both hemispheres returns to the pre-eruption value after 15–20 years (Fig. S12c,d). However, Zhong et al. (2011) find that these sea ice changes following volcanic eruptions depend on the initial state. Centennial sea ice expansion that sustains cold anomalies could thus be important for some eruptions and could have played a role in initiating the LIA (Miller et al., 2012).
5.5 Consistency with simulations
CMIP6 last-millennium simulations disagree markedly among themselves and with our reconstruction. They do not capture the millennial-scale cooling trend (Fig. 8) and the OHC loss after a series of volcanic eruptions differs much in magnitude (Fig. S15). While the lack of a global-mean cooling trend, or even the presence of strong warming, is the main reason for the mismatch in the other fields, they also disagree in the spatial structure, particularly in the mid and high latitudes. While proxy reconstructions have their own issues (Anchukaitis and Smerdon, 2022; Zhu et al., 2020), this potentially highlights the value of DA, which adjusts the model state through time using proxy information.
@@ 5.5 Consistency with simulations
Since the annual-mean cooling trend is likely volcanically forced (Schurer et al., 2014; Büntgen et al., 2020; Wanner et al., 2022), the lack of a trend may imply that the coupled climate system response to volcanic eruptions, such as post-eruption OHC loss, is not well resolved in climate models (Church et al., 2005). Other model issues that can lead to discrepancies with proxy reconstructions include uncertainty in the forcing dataset and in the aerosol representation (Lücke et al., 2023; Marshall et al., 2022; Timmreck, 2012). A false warming trend can also arise from impulsive positive radiative forcing when the prescribed background aerosols from the spin-up run are removed at the start of a last-millennium run (Gregory et al., 2013; Fyfe et al., 2021).
@@ Acknowledgements
We thank Zilu Meng, Eric J. Mei, Kyle C. Armour, and Vincent T. Cooper for helpful conversations. Anna Black’s master’s thesis provided preliminary results that helped motivate this study. This material is based upon work supported by the National Science Foundation under Award No. 2202526. We would like to acknowledge computing support from the Casper system (https://ncar.pub/casper) provided by the NSF National Center for Atmospheric Research (NCAR), sponsored by the National Science Foundation. This work made use of pyleoclim (Khider et al., 2022) and climlab (E. J. Rose, 2018).
Data availability
The reconstruction data and code will be published once this manuscript has been accepted.

### v2

@@ 5.4 Climate response to volcanic eruptions
Large volcanic eruptions are the dominant pre-industrial climate forcing (Büntgen et al., 2020). Clusters of them are thought to be the main cause of the LIA (Miller et al., 2012; Brönnimann et al., 2019; Gupta and Marshall, 2018) and previous cold periods (Van Dijk et al., 2024). Sea-surface cooling after eruptions is initially damped by the upper ocean and absorbed into the deeper ocean (Gregory et al., 2016). The cold anomalies at depth re-emerge over subsequent years, extending surface cooling well beyond the volcanic aerosol residence time of 1–2 years. In our volcanic composite, OHC does not recover until 10–20 years after the eruption (Fig. 10b). A decadal pacing of large eruptions would thus allow heat loss to accumulate and persist (Gupta and Marshall, 2018; Zhong et al., 2011). Our results show OHC anomalies of up to $80\,\mathrm{ZJ}$ for volcanic clusters during the late 1200s, late 1400s, early 1700s, and early 1800s (Fig. S17). Miller et al. (2012) hypothesize that the late 1200s cluster (annotated in Fig. 5b), with major eruptions in 1276 and 1286, initiated the cooling trend leading to the LIA. Indeed, the 20-yr period with the lowest EEI over the last millennium is 1272–1291 ($0.4\,\mathrm{W\,m^{-2}}$), with a minimum in 1284 ($0.7\,\mathrm{W\,m^{-2}}$). This is decades after the 1257 Samalas eruption, which had an exceptional magnitude but a relatively weak climate response (Fig. S17a; Vidal et al., 2016; Guillet et al., 2017). The 20-yr period with the largest loss of OHC300 is 1803–1822 ($0.4\,\mathrm{W\,m^{-2}}$; Fig. S17d), including the 1815 Tambora eruption, and the recovery from this cluster marks the end of the LIA (Brönnimann et al., 2019). This early 1800s cluster complicates the definition of a pre-industrial baseline (e.g., Thorne et al., 2026) and the initialization of historical simulations in 1850 (Ballinger et al., 2026).
@@
OHC anomalies in our reconstruction differ from model simulations (Figs. 10, S16, and S17). Simulated ocean cooling peaks 3 years after the eruption and recovers immediately, while the reconstructed cooling peaks at a lower amplitude. In contrast to the simulations, the reconstructed onset of recovery varies by eruption, occurring 3–10 years post-eruption and resulting in the flatter composite (Fig. 10b). The reconstructed response is likely smoothed by biological memory, or persistence, in TRW proxies, which form the bulk of our network (Fritts, 1976; Esper et al., 2015; Lücke et al., 2019; Zhu et al., 2020). This results in a muted and time-lagged volcanic response, which has long been noted in the literature (e.g., Frank et al., 2007; Anchukaitis et al., 2012; Mann et al., 2012; D’Arrigo et al., 2013; Stoffel et al., 2015; Anchukaitis and Smerdon, 2022; He et al., 2026). However, models are also known to overestimate the cooling after large-magnitude eruptions, for example, due to the way volcanic aerosols are prescribed and volcanic forcing is calculated (Timmreck et al., 2009; Timmreck, 2012; Marotzke and Forster, 2015; LeGrande et al., 2016; Marshall et al., 2025). The model–proxy discrepancy is further exacerbated by unforced variability and uncertainty in VSSI and timing (Zanchettin et al., 2019; Lücke et al., 2023), although progress has been made in resolving large eruptions in ice cores (Burke et al., 2019) and reconciling simulated and reconstructed volcanism (Stoffel et al., 2015; Zhu et al., 2020). Over 1960–2000, the timescale and magnitude of OHC loss in our reconstruction agree with instrumental datasets for three large eruptions (Fig. 4).
The loss of OHC is driven by a negative EEI at the TOA, peaking at $0.2\,\mathrm{W\,m^{-2}}$ in the composite, and arising from increased RSR (Fig. 10c,d). This is much weaker than expected for eruptions larger than Mt. Pinatubo in 1991, which had a peak EEI of around $3\,\mathrm{W\,m^{-2}}$ (Allan et al., 2014). After 3–5 years, the EEI becomes marginally positive due to reduced RSR, which is also evident in coupled model simulations, although at much larger magnitudes (Fig. S16c). By Eq. (1), the difference between dOHC300/dt $\approx C(dT/dt)$ and the EEI may be due to heat uptake $\gamma T$ by the deeper ocean. However, the difference here is likely due to the greatly underestimated volcanic forcing. The ocean cooling dOHC300/dt, with a peak of $0.4\,\mathrm{W\,m^{-2}}$ in the composite, may be a better estimate of the volcanic forcing than the reconstructed EEI.
@@
Previous modeling studies have found expanded sea ice persisting for more than a century following large eruptions due to sea-ice–ocean feedbacks in the Arctic Ocean and North Atlantic (Zhong et al., 2011; Miller et al., 2012; Slawinska and Robock, 2018). In contrast, in our volcanic composite, the sea-ice area in both hemispheres returns to the pre-eruption value after 15–20 years (Fig. S15c,d). However, Zhong et al. (2011) find that these sea ice changes following volcanic eruptions depend on the initial condition of the coupled atmosphere–ocean–sea-ice system. Centennial sea-ice expansion that sustains cold anomalies could thus be important for some eruptions and could have played a role in initiating the LIA (Miller et al., 2012).
5.5 Comparison with last-millennium simulations
The majority of CMIP6 models do not show the millennial-scale cooling trend found in proxy reconstructions (Fig. 8 and Section 44.4; PAGES 2k Consortium, 2019), in contrast to CMIP5 models (Atwood et al., 2016). The lack of a global-mean cooling trend over 850–1850 in models is the main reason for the difference in OHC and TOA radiation with our reconstruction. If we assume that trends of all fields scale linearly with the global-mean temperature trend and adjust the model trends to match the reconstructed temperature trend, agreement improves substantially, particularly for the global means (not shown). The benefit of DA is thus that variability on seasonal to millennial timescales is directly constrained to follow the climate trajectory as sampled by proxies. Model biases may be inherited by the LIMs and thus the reconstruction, which is partly mediated by our multi-model reconstruction.
@@ 5.5 Comparison with last-millennium simulations
Our reconstruction provides a new dataset to compare with energy budgets simulated in coupled models. First, since the annual-mean cooling trend is likely volcanically forced (Schurer et al., 2014; Büntgen et al., 2020; Wanner et al., 2022), the lack of a simulated trend may imply that the coupled climate system response to volcanic eruptions, such as post-eruption OHC loss, is not well resolved in current models (e.g., Church et al., 2005). Other model issues that can lead to discrepancies with proxy reconstructions include uncertainty in the forcing dataset and in the aerosol representation (Lücke et al., 2023; Marshall et al., 2022; Timmreck, 2012). A false warming trend can also arise from impulsive positive radiative forcing when the prescribed background aerosols from the spin-up run are removed at the start of a transient last-millennium simulation (Gregory et al., 2013; Fyfe et al., 2021). Finally, the initialization of coupled model simulations from equilibrium is likely a poor approximation of the coupled climate system. This critique applies to both historical simulations starting in 1850, which are preceded by strong volcanism and the last-millennium cooling trend (Ballinger et al., 2026; Gebbie and Huybers, 2019), and to last-millennium simulations starting in 850, when the deep ocean is potentially still adjusting to the last deglaciation (Zhu et al., 2019). Proxy-constrained estimates of the pre-industrial deep ocean may help to provide initial conditions that better reflect earlier climate variability.
6 Conclusion
Earth’s energy imbalance is a fundamental climate metric, but our understanding of its variability on long timescales is limited by a lack of long records and large uncertainty in coupled climate models. Our reconstruction over the last millennium extends the observational record by 1000 years, providing context for energy budget variability in the absence of anthropogenic forcing. By combining proxy data with a simplified climate model using data assimilation, we obtain a consistent dataset of global surface temperatures, sea ice, top-of-atmosphere radiation, and upper-ocean heat content. In contrast to climate models, the variability on all timescales is directly constrained by proxies to follow the observed climate trajectory. The reconstruction can be used to investigate seasonal to multidecadal variability, millennial trends, and episodic events like volcanic eruptions.
We demonstrate the feasibility of reconstructing top-of-atmosphere radiation from temperature-sensitive proxies, even though the quantification of the energy imbalance and ocean heat content remains challenging during the satellite period (Hakuba et al., 2024). Through validation in pseudoproxy experiments and against satellite observations, we find that the global-mean energy imbalance, reflected shortwave radiation, and outgoing longwave radiation can be skillfully reconstructed; there is less skill in spatial patterns, especially outside of the tropics. Volcanic forcing events and a subsequent reduction in ocean heat content are captured, but with greatly underestimated magnitude. Since the simplified climate model we use is trained on last-millennium simulations, our reconstruction resolves natural forcing and the radiative response, but not anthropogenic forcing. Compared to the reconstruction of the last millennium by Gebbie and Huybers (2019), we find agreement on the magnitude of sustained energy loss, but with large error bars around zero and uncorrelated multidecadal variability, likely due to different proxy networks and reconstruction methods. These uncertainties could potentially be be reduced by including constraints on the deep ocean from proxies, either in situ (e.g., Waelbroeck et al., 2002; Lear et al., 2000) or from globally integrating noble gases (Bereiter et al., 2018a). These proxies have proven useful for reconstructions of the last deglaciation (Bereiter et al., 2018b; Baggenstos et al., 2019; Shackleton et al., 2023) and the Pleistocene (Shackleton et al., 2026).
Our reconstruction provides context for the energy gain in recent decades, which we find to be unprecedented in the context of natural, pre-industrial variability. It also reaffirms the role of volcanism in the last-millennium cooling trend, which may have been continuously forced by volcanic clusters causing an accumulation of heat loss. The discrepancy of our reconstruction with coupled simulations may point to deficiencies in the energy budget representation in the same models used for future climate projections.
@@ Acknowledgements
We thank Jess Tierney, Sylvia Dee, and an anonymous reviewer for their comments in review, which greatly improved this manuscript. We thank Kyle Armour, Zilu Meng, Eric Mei, Aaron Donohoe, and Vince Cooper for helpful conversations. Anna Black’s master’s thesis provided preliminary results that helped motivate this study. This material is based upon work supported by the National Science Foundation under Award No. 2202526. We would like to acknowledge computing support from the Casper system (https://ncar.pub/casper) provided by the NSF National Center for Atmospheric Research (NCAR), sponsored by the National Science Foundation. This work made use of pyleoclim (Khider et al., 2022) and climlab (E. J. Rose, 2018).
Code and data availability
Code is publicly available at https://github.com/DominikStiller/energy-budget-from-proxies. The reconstruction data will be published once this manuscript has been accepted.